x-3 square root of x -4=0
I could not find the square root symbol on my keyboard

Answer by stanbon(75887)

x-3 square root of x -4=0
----
x - 3sqrt(x) - 4 = 0
----
3sqrt(x) = x-4
----
Square both sides to get:
9x = x^2 - 8x + 16
----
x^2 -17x + 16 = 0
Factor:
(x-1)(x-16) = 0
x = 1 or x = 16
------
Check for extraneous solutions:
x - 3sqrt(x) - 4 = 0
Test x= 1
---
1 - 3 - 4 = 0 :::: false
-----
Test x = 16
16 - 3*4 - 4 = 0
16 - 16 = 0 ::::: true
----
Colclusion: x = 16
======================
Cheers,
Stan H.